OpenStudy (anonymous):

A hotel has rooms with one or two beds. Total 19 rooms and 31 beds. If e is the rooms with one bed and z is the rooms with two beds, then e-z= ?

OpenStudy (anonymous):

its a mulple choice question...a)5 b)14 c)-14d) 24 e)-5 if that helps?

OpenStudy (anonymous):

\[e+z=19\] \[e+2z=31\] so \[z=12\] and therefore \[e=7\]

OpenStudy (anonymous):

finally \[e-z=7-12=-5\]

OpenStudy (anonymous):

awesome! thank you! i wish i listened to my teacher blahbinn more